At early stance phase, a runner’s ground reaction force is 1500 N at a 55-degree angle with respect to the ground. If the runner’s mass is 90 kg, what are their accelerations ax and ay?
Answer #1

Ground reaction forces are usually upwards.

The statement can be made into a diagram which will look something like this

1500 Ground

so,

we can resolve the force into components

Fx = 1500 * cos55 = 860.36 N

Fy = 1500 * sin 55 = 1228.72 N

so

Fx = max

ax = Fx / m

ax = 9.56 m/s2

and

ay = Fy / m

ay = 13.65 m/s2